### v3

# A Survey of Fuzzing Open-Source Operating Systems

Abstract.
Vulnerabilities in open-source operating systems (OSs) pose substantial security risks to software systems, making their detection crucial. While fuzzing has been an effective vulnerability detection technique in various domains, OS fuzzing (OSF) faces unique challenges due to OS complexity and multi-layered interaction, and has not been comprehensively reviewed. Therefore, this work systematically surveys the state-of-the-art OSF techniques, categorizes them based on the general fuzzing process, and investigates challenges specific to kernel, file system, driver, and hypervisor fuzzing. Finally, future research directions for OSF are discussed.
Operating System Fuzzing, Kernel Fuzzing, File System Fuzzing, Driver Fuzzing, Hypervisor Fuzzing
copyright: acmlicensedccs: Security and privacy Systems securityccs: Software and its engineering Software testing and debugging
1. Introduction
Operating systems (OSs) play a crucial role in the operation of modern computer systems. They are responsible for managing the hardware and software resources of a computer. They serve as the cornerstone for maximizing hardware resource utilization and ensuring software system stability. Specifically, open-source OSs, such as Linux, Android, FreeRTOS, Zephyr, etc., have become particularly favored in most fields like autonomous driving, robotics, cloud services, and Web of Things for the advantages of transparency, customizability, and community-driven innovation, leading to their anticipated dominance in future computer systems.
However, as the scale of open-source OS codebases continues to expand, the security threats they pose have become increasingly alarming, raising significant concerns about the secure operation of software systems. According to National Vulnerability Database (NVD) (NVD), more than 9,300 vulnerabilities have been discovered in the kernels of mainstream open-source OSs (including Linux, Android, FreeBSD, OpenBSD and Zephyr) since 2004. In particular, the number of vulnerabilities surges to 3,300 in the single year of 2024, an increase that is 10 times higher than the previous year. Similarly, the Common Vulnerabilities and Exposures (CVE) (cve) has documented at least 7,600 vulnerabilities in the Linux kernel to date. Among them, NVD has disclosed at least 1,566 high-severity vulnerabilities and 156 critical vulnerabilities that require immediate remediation.
OS-level vulnerabilities are particularly concerning when compared to those in user-level applications. This is because such vulnerabilities can be further exploited, potentially allowing attackers to gain complete control over the OSs (Xu2015FromCT; Zhang2015AndroidRA; chen2019slake), leading to irreversible consequences. The staggering number of these vulnerabilities and the malicious outcomes they have caused have attracted significant attention from security researchers. As a result, there is a strong and growing interest in developing effective and efficient techniques to identify and mitigate these potential vulnerabilities, thereby aiding the continuous evolution of open-source OSs. One of the widely used techniques is fuzzing, which was first introduced by Miller et al. (Miller1990AnES) in 1990, and has achieved notable success across various domains, e.g., compilers (Yang2023WhiteboxCF), interpreters (Holler2012FuzzingWC), and open-source software (Serebryany2017OSSFuzzG).
Fuzzing, also known as fuzz testing, is a technique that involves feeding semi-randomly generated test cases as inputs to the program under test (PUT) to trigger program paths that may contain software vulnerabilities. Over the past decade, fuzzing has become capable of effectively testing complex OS code. This progress has received widespread attention from researchers, who aim to enhance the depth and breadth of OS fuzzing (OSF) by incorporating cutting-edge techniques such as program analysis and deep learning. However, compared to traditional fuzzing, the complex domain knowledge involved in OS makes developing effective and efficient OSF particularly challenging. It not only requires focusing on advancements in fuzzing techniques, but also demands consideration of the inherent complexity and multi-layered interaction of OS. Typically, survey papers play a key role in advancing this field by providing a comprehensive review of the OSF techniques as well as summarizing existing challenges and pinpointing potential directions.
However, to date, there has been no systematic review of OSF. While there are some survey papers on traditional software fuzzing (Eisele2022EmbeddedFA; Yun2022FuzzingOE; Zhu2022FuzzingAS; Mans2018TheAS; Li2018FuzzingAS; Liang2018FuzzingSO; Zhang2018SurveyOD; Mallissery2023DemystifyTF; bohme2020fuzzing; Godefroid2020), they do not systematically introduce the general steps of OSF, and they also do not highlight the unique challenges that OSF faces compared to traditional software fuzzing. Therefore, to bridge the gap, a systematic review of the state-of-the-art OSF is essential, aiming to provide a comprehensive guideline on this topic.
For the scope of this survey, after conducting a systematic search of state-of-the-art OSF, we observed that the PUT is typically classified into kernel, file system, driver, and hypervisor because each of these OS layers present distinct challenges in fuzzing. Specifically, in a highly heterogeneous hardware environment (e.g., for intelligent vehicles), achieving unified resource management within the OS requires the support from hypervisor; and vulnerabilities from hypervisor can be exploited to launch malicious attacks against the host OS. Therefore, to ensure the completeness of our survey, we also included hypervisor, as a component of a generalized OS, within our scope.
To systematically review OSF, we designed a comprehensive search process to identify existing high-quality research materials (see Section 2), including 60 state-of-the-art OSF papers and 4 open-source OSF tools. Through a thorough analysis of these materials, we uncovered the rising trend in OSF research and explained the reasons behind it. Then, we introduced the distinctive features of fuzzing techniques for the four OS layers (i.e., kernel, file system, driver, and hypervisor) from a high-level perspective, and summarized a general workflow of OSF, consisting of three core modules, i.e., input, fuzzing engine, and running environment (see Section 3). Next, we conducted a comprehensive review of the state-of-the-art OSF with respect to seven key steps in the three core modules (see Section 4), aiming to systematically sort out the technological advancements in OSF and the associated complex domain knowledge. Meanwhile, we summarized the unique problems encountered by the four OS layers during fuzzing as well as the corresponding solutions (see Section LABEL:Section5). Finally, based on our findings, we provided four future research directions in the OSF field (see Section LABEL:Section7).
2. Collection Strategy and Result
Following (garousi2016systematic), we use a scientific and effective collection strategy (see Section 2.1) and present a detailed analysis of the collection result (see Section 2.2) to systematically review OSF.
2.1. Collection Strategy
Figure 1 shows our strategy to collect relevant works, assess their quality, and keep updated with the latest publications. We both review scientific literature and collect open-source tools for OSF.
2.1.1. Scientific Literature Review
We set the temporal scope of our survey to cover the period from the earliest relevant papers or tools in this field to the present, ranging from January 2015 to August 2024 (i.e., around 10 years). The detailed steps are as follows.
Figure 1. Our Collection Strategy
Database Search. This step aims to find the potential relevant papers by searching electronic databases. Specifically, we select ACM Digital Library, IEEE Xplore, DBLP and Semantic Scholar as our databases, which are popular bibliography databases containing a comprehensive list of research venues in computer science. Initially, employing “Operating System Fuzzing” as the sole keyword fails to fully capture the existing literature about OSF. Keywords need to focus on the application of fuzzing to operating systems (e.g., Linux, Android, FreeBSD, etc.). In addition, keywords should cover the main tested objects in open-source operating systems, such as kernel, file system, driver, etc. We optimize the search keywords in an iterative manner for the purpose of collecting as many related papers as possible. Our final search keywords are reported as follows. Moreover, our search targets titles, abstracts, and keywords of the papers, since these parts often convey the theme of a paper. Finally, we obtain a total of 56 candidate papers during database search.
(“Linux” OR “Android” OR “FreeBSD” OR “OpenBSD” OR “Zephyr” OR “Open-source Operating System”)
AND (“Kernel” OR “File system” OR “Driver” OR “Hypervisor”) AND (“Fuzzing” OR “OSF”)
Paper Filtering. We perform a manual assessment on the 56 candidate papers obtained from our database search to ensure their relevance and quality. Specifically, to determine whether each candidate paper is relevant to OSF and has high quality, we analyze the abstracts and introductions of these papers, following the inclusion and exclusion criteria formulated as follows. We use two inclusion criteria, i.e., IC1: papers that introduce the process of OSF, and IC2: papers that propose a technique of OSF. We use four exclusion criteria, i.e., EC1: survey papers or summary papers, EC2: papers that do not target fuzzing, EC3: papers that do not focus on fuzzing operating system and its components, and EC4: papers that have not been published in top-tier conferences or journals. Specifically, for EC1, such survey papers are discussed in Section 3.1 for a comparison with our survey; and for EC4, we discuss these top-tier conferences and journals in Section 2.2. Through our manual assessment, we remove 13 papers, resulting in 43 papers.
Backward & Forward Snowballing. To reduce the risk of missing relevant papers, we perform both backward and forward snowballing (wohlin2014guidelines) on the 43 papers. In backward snowballing, we check the references in these papers to obtain candidate papers, while in forward snowballing, we use Google Scholar to locate candidate papers that cite these papers. For these candidate papers obtained by snowballing, we also apply the same inclusion and exclusion criteria to identify relevant papers. Finally, we add 17 new relevant papers, resulting in a final set of 60 papers.
Full-Text Analysis. We download all the resulting 60 papers, and conduct a full-text analysis to identify the fuzzing target (i.e., the operating system and its components) and the proposed fuzzing technique. After reading all these papers, we classify them to form our survey (see Section 4 and LABEL:Section5).
2.1.2. Open-Source Tools Collection
Some open-source tools of OSF have not been published in academic papers, and these tools should not be ignored. Therefore, we use the same search keywords to collect open-source tools whose stars are more than 400 stars on GitHub. We eliminate tools that have been published in academic papers, and select 4 additional open-source tools.
(a) Distribution across Publication Venues
(b) Distribution Across Publication Years and Layers
Figure 2. The Analysis Results of Paper Collection
2.2. Collection Result Analysis
We analyze the collected papers from three perspectives, i.e., the publication venues, the publication years, and the target OS layers.
Publication Venues. Figure 2(a) shows the distribution of all papers across the publication venues. The 60 papers are published across 18 top-tier venues in four domains, i.e., security, software engineering, computer architecture, and computer storage systems. Specifically, (i) most of the papers, up to 77%, are published in security venues such as USENIX Security Symposium (USEC), ACM Conference on Computer and Communications Security (CCS), IEEE Symposium on Security and Privacy (S&P), and Network and Distributed System Security Symposium (NDSS); (ii) 11% are published in software engineering venues such as ACM Symposium on Operating Systems Principles (SOSP), International Symposium on Software Testing and Analysis (ISSTA), and International Conference on Software Engineering (ICSE); (iii) there are 5 papers, accounting for 10%, published in computer architecture venues, one each in European Conference on Computer Systems (EuroSys), USENIX Annual Technical Conference (ATC), ACM Transactions on Embedded Computing Systems (TECS), IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems (TCAD), and International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS); and (iv) since operating system is related to storage system, there is 1 paper published in computer storage system venues (i.e., ACM Transactions on Storage (TOS)). It can be conclude that application of fuzzing techniques to operating systems spans multiple fields of computer science.
Publication Years and Target OS Layers. Figure 2(b) presents the number of papers published in each year as well as the distribution across the target OS layers in each year. Overall, the number of OSF papers shows a general ascending trend from 2015 to 2024. It is evident that interest in OSF has been escalating year by year, which indicates increased attention in leveraging fuzzing to uncover deeper security vulnerabilities in open-source OS. Notice that since the papers from 2024 have not been fully surveyed yet, Figure 2(b) only shows the number of OSF publications till August 2024. Moreover, most papers focus on the fuzzing of kernel. Particularly, Google’s fuzzing framework Syzkaller (Syzkaller) launched in 2015. It provides researchers with a foundational kernel fuzzing engine, and has made a substantial impact on kernel fuzzing. However, using kernel fuzzing interfaces (i.e., between user and kernel space) fails to uncover deep security vulnerabilities in other OS layers (e.g., verification chain checks in driver). Therefore, an increasing number of papers realized the necessity of fuzzing other OS layers (i.e., file systems, drivers, and hypervisors) since 2020.
3. Overview of OSF
After analyzing existing fuzzing surveys, we employ PUT-based classification for a systematic review of OSF (Section 3.1). Following this classification, we introduce the main tasks of the four OS layer fuzzing (Section 3.2), and provide a high-level overview of the general workflow of OSF (Section 3.3).
3.1. Classification Dimension
Existing fuzzing surveys classify the literature by one of the four dimensions, i.e., 1) the amount of information the fuzzer requires or uses (black-, white-, and gray-box) (Mans2018TheAS; Li2018FuzzingAS; Liang2018FuzzingSO; Godefroid2020; bohme2020fuzzing), 2) the strategy employed for seed update (mutation- and generation-based) (Saavedra2019ARO; schumilo2020hyper; schumilo2021nyx; myung2022mundofuzz; pan2021V-shuttle; cesarano2023iris; bulekov2022morphuzz), 3) the research gaps of integrating advanced techniques into traditional fuzzing workflows (Wang2019ASR; Zhang2018SurveyOD; Wang2020SoKTP; Zhu2022FuzzingAS; Mallissery2023DemystifyTF), and 4) the PUT (Yun2022FuzzingOE; Eisele2022EmbeddedFA). Here, we adopt the fourth dimension as the guiding taxonomy, as the first three dimensions primarily focus on the general study of fuzzing techniques and lack the specificity required for analyzing OSF. In contrast, the PUT-based classification provides a comprehensive framework for systematically reviewing the general workflow of OSF while focusing on the unique challenges posed by the different OS layers (i.e., different PUTs).
3.2. OS Layers
According to the PUT-based classification we employed, OSF can be categorized into four types, i.e., kernel fuzzing, file system fuzzing, driver fuzzing, and hypervisor fuzzing. Due to the distinct differences in fuzzing approaches across these four OS layers, it is necessary to outline the primary functions of each OS layer and the primary tasks of their respective fuzzing approaches.
(a) Interfaces of Each OS Layer
(b) General Workflow of OSF
Figure 3. Interfaces of Each OS Layer (where 1, 2, 3 and 4 denote the fuzzing interfaces for Kernel, Driver, File System and Hypervisor respectively; and 1 is also the fuzzing interface for Driver and File System); and General Workflow of OSF (which has three modules, i.e., input, fuzzing engine, and running environment)
Kernel. Kernel is one of the most critical systems in an OS because it manages essential resources such as processes and memory for the entire system and provides a unified programming interface for user-space programs to interact with hardware resources. All other system types rely on and operate on top of the kernel. Consequently, vulnerabilities in the kernel can be maliciously exploited, potentially causing significant damage to the systems running on it. As illustrated in Figure 3(a), to perform kernel fuzzing, a fuzzer triggers the kernel’s code paths by switching from user space to kernel space through syscall 1 (a.k.a. system call) (Koopman1997ComparingOS). Therefore, the core task of a kernel fuzzer is to generate various test cases (referred to as seeds) by combining syscalls provided by the kernel to continuously trigger the kernel’s code paths. For example, the darker sequences in Figure 3(a) denote branches covered by test cases, while the lighter circles denote code blocks yet to be covered.
File System. As a core system service within an OS, file system is essential for tasks such as reading, writing, managing, and scheduling files, as well as ensuring data consistency during system crashes. Most file systems, like ext4 (Cao2007Ext4TN), XFS (XFS2018), Btrfs (Rodeh2013BTRFSTL) and F2FS (Lee2015F2FSAN), operate within the OS kernel. Therefore, approaches used in kernel fuzzing can be adapted for file system fuzzing. However, approaches based purely on syscalls often yield numerous invalid results because the system state is predominantly influenced by metadata in file system operations. In contrast, operations on regular file data through syscalls like read() and write() contribute little to identifying file system vulnerabilities. Hence, an effective file system fuzzer typically combines sequences of file operation-related syscalls 1 with images 3 where metadata has been changed. As shown in Figure 3(a), the altered disk image is mounted to the file system’s partition using privileged commands (schumilo2017kafl; Hydra2020finding), becoming the new target for fuzzing. Kernel-provided syscalls are then used to conduct read, write, management, and scheduling operations, facilitating a comprehensive fuzzing of the file system.
Driver. Drivers are responsible for communication and control between the user/kernel and hardware devices. They act as the bridge between hardware and OS, managing tasks such as device initialization, data transfer, and interrupt handling, ensuring that user-space applications can interact with underlying hardware (Chou2001AnES; Palix2011FaultsIL; Stoep2018AndroidSecurity). In driver fuzzing, the primary objective is to uncover vulnerabilities in device drivers across the initialization, data communication, and control stages. Since drivers can receive operation requests from both user space and hardware devices, they expose a broader attack surface compared to the kernel or other kernel subsystems (Beniamini2017Part1; Beniamini2017Part2; Chang2017; Davis2011; NohlLell2014). As illustrated in Figure 3(a), the two main attack surfaces for driver fuzzing are syscalls 1 and peripheral interfaces 2. Unlike kernel fuzzers, syscall-based driver fuzzers focus more on syscalls that directly operate on device files, e.g., read(), write(), seek(), ioctl(), etc.
The peripheral interface can be exploited in two ways, i.e., I/O interception and device configuration. I/O interception involves intercepting access to I/O objects (e.g., DMA, MMIO, and Port I/O) to mutate I/O data, which is then redirected to the target to observe the driver’s behavior. Device configuration, on the other hand, simulates peripheral device behavior and injects the simulated data into the I/O channel. In addition, it is worth noting that drivers constitute the largest codebase among kernel subsystems, and exhibit significant variability due to implementations by different vendors. As a result, driver fuzzers often face larger challenges in achieving high coverage, generating diverse test cases, and ensuring fuzzing effectiveness given these characteristics.
Hypervisor. In environments with highly heterogeneous hardware, OSs inherently lack the capability to manage and schedule heterogeneous computing resources, such as those used in industries like railways, avionics, and automotive systems (Cinque2021VirtualizingMS). To enable the concurrent execution of multiple OSs while maintaining secure isolation between them in such heterogeneous resource environments (RTCA_DO_178C; ISO_26262_2011), hypervisors, a.k.a. Virtual Machine Monitors (VMMs), leverage virtualization technique to partition hardware resources (e.g., CPU, memory, and disk space) into multiple virtual partitions (Popek1974FormalRF; Cinque2021VirtualizingMS; Cilardo2021VirtualizationOM). When a guest OS, which runs within a Virtual Machine (VM), accesses or operates hardware devices, it triggers a VM-exit event that transfers these privileged operations to the hypervisor for hardware emulation. Therefore, the objective of hypervisor fuzzing is to accurately emulate the behavior of the guest OS when accessing and operating these virtualized hardware components, which serves as the primary entry point for implementing hypervisor fuzzing trials. As illustrated in Figure 3(a), the main interfaces involved include I/O channels and hypercalls 4.
To ease the understanding of how to perform hypervisor fuzzing using these interfaces, we introduce the technical background of hypervisors. Technically, hypervisors can be categorized into full virtualization and para-virtualization. In full virtualization, the guest OS is unaware of its virtualized environment. Consequently, when the guest OS attempts to access physical hardware (e.g., memory access), the hypervisor intercepts the request. At this point, the VM triggers a trap and enters a VM-exit state, with the hypervisor assuming a full control of the VM’s operations. Subsequently, the hypervisor forwards the memory access request to the Device Emulator, which provides virtualization for Port I/O (input/output port instructions), MMIO (Memory-Mapped I/O for direct memory access), and DMA (for complex and large-scale data transfers). Finally, the Device Emulator returns the virtual memory access interface to the guest OS. In a fully virtualized environment, the I/O channel is more extensively utilized because this technique offers complete hardware emulation. However, this can also lead to more frequent communication overhead.
To facilitate and accelerate communication between the guest OS and the hypervisor, modern hypervisors often support hardware-accelerated virtualization technologies. By introducing specialized instructions (such as vmcall) to perform hypercalls, para-virtualization allows OSs to bypass the VM and communicate directly with the hypervisor. Therefore, in addition to the I/O channel, the hypercall interface can also be employed to implement hypervisor fuzzing trials. As a result, it is necessary to conduct specialized fuzzing of hypervisors to uncover vulnerabilities related to resource management and security isolation in generalized OSs.
3.3. OSF Workflow
We summarize the general workflow of OSF in Figure 3(b), which consists of three key modules, i.e., input, fuzzing engine, and running environment. Each module is described in detail as follows.
Input. There are three types of inputs to be considered in the input module, i.e., initial seed, specification, and target OS. First, the initial seed is the raw material fed to the fuzzing engine, which is initially deposited in a global corpus (i.e., a repository for storing seed candidates) and subsequently used throughout the fuzzing workflow. Second, the specification, which describes the structure and syntax of the seed, is parsed by the fuzzing engine and used to generate new seeds or enhance the quality of the seeds. It is important to note that the use of the specifications is not applicable to all scenarios; it mainly applies to highly structured seeds with documented descriptions of the seed structure (Syzkaller; chen2020koobe; zhao2022statefuzz; sun2021healer; shen2021rtkaller; shen2022tardis; jeong2023segfuzz; xu2020krace; jeong2019razzer). Third, the target OS requires to be instrumented to collect runtime information for feedback analysis and bug monitoring during fuzzing.
Fuzzing Engine. The main function of the fuzzing engine is to generate new seeds and wrap them into test cases for the input fed to the executor in the running environment. This module go through a complete closed loop of seed selection from a corpus, seed mutation and/or generation, and seed trim before storing to the corpus. Specifically, the corpus stores only seeds that have been verified as “high-quality”, where “high-quality” seeds can be defined as the seeds that have triggered bugs (zou2022syzscope; wu2018fuze; lin2022grebe; shen2021rtkaller), the initial seed generated based on tailored rules (shen2022drifuzz) or specifications (Syzkaller), or the streamlined seeds after seed trim as they trigger new coverage (pailoor2018moonshine; shen2021rtkaller). Such a consideration stems from the empirical conclusion that mutations based on high-quality seeds usually have more opportunities to move the execution of the PUT closer to trigger bugs (you2017semfuzz; shen2022drifuzz; lin2022grebe; Unicorefuzz2019). However, even if the corpus always stores high-quality seeds, it is necessary to consider which seeds should be prioritized in order to improve the efficiency of seed mutation. Therefore, seed selection is often adopted to prioritize the seeds in the corpus, and a computation method of seed priority is often devised to ensure that the selected seeds have more chances to reveal new bugs, e.g., PageRank (lin2022grebe), program analysis (wu2018fuze; chen2022sfuzz), evolutionary algorithm (shen2022drifuzz; henderson2017vdf; xu2020krace), empirical strategy (xu2020krace; Syzkaller; you2017semfuzz; zou2022syzscope), reinforcement learning (wang2021syzvegas), etc. Then, to continuously generate test cases to automate the whole fuzzing process, seed mutation and/or generation transform and/or generate the seeds in some way and wrap them into test cases that can be executed by the target OS. For example, seed mutation employs parameter-level bit flips or byte-level replacement to generate seeds based on the high-quality seeds in the corpus. Seed generation builds test case models from the specification to generate new seeds. Finally, to further improve the speed performance of OSF, the generated seed usually requires to be trimmed to reduce its mutation space as well as its execution time, e.g., greedy algorithm (pailoor2018moonshine) and stepwise filtering (schumilo2020hyper; henderson2017vdf). The trimmed seeds are then stored into the corpus, and the process is repeated to automatically drive new and high-quality seed generation.
Running Environment. The running environment is responsible for feeding the test cases wrapped by the fuzzing engine into the executor. Unlike traditional fuzzing techniques, OSF usually relies on full-virtual environment to avoid as much as possible the effects propagated by OS crashes. In contrast, traditional fuzzing techniques do not fatally affect the entire fuzzing process even if a crash occurs. Therefore, the running environment of OSF is often designed as full-virtual environment that can be quickly recovered, allowing the entire system to be quickly restarted to a known, clean state in the event of a crash (shen2022tardis; lin2022grebe; schumilo2021nyx; pan2021V-shuttle; zhao2022semantic; song2020agamotto). Such a design allows a fuzzer to run thousands of test cases continuously and automatically in an isolated environment without worrying about the potential for lasting damage from individual test cases. Subsequent execution results fall into two categories. If the executor does not crash, the feedback will be collected by the instrumented code, which is used to guide seed mutation of the fuzzing engine to steer the fuzzing towards uncovered program paths. If the executor triggers a crash, the instrumented code records the information and sends it to the bug analyzer. The bug analyzer identifies and categorizes errors, and generates detailed bug reports. Finally, if the fuzzing process is not yet finished, it should be restored to its pre-crash state, and then move on to explore other states of the target OS.
Although there are some similarities with traditional fuzzing in terms of the coarse-grained fuzzing process, the technical challenges and implementation details of each step need to take into account the specificity and complexity of OS, which ultimately makes it show a big difference with traditional fuzzing in terms of the fine-grained details. Therefore, we will elaborate on the technical details and technical distinctions of each step illustrated in Figure 3(b) in Section 4 and LABEL:Section5.
4. General Workflow Characteristics of OSF
We review the state-of-the-art OSF with respect to the key steps in the three core modules of the general fuzzing workflow, i.e., input, fuzzing engine, and runtime environment, as shown in Figure 3(b). Our review results are reported in Table LABEL:tab1, where “-” means that this dimension is not relevant, or not clearly mentioned in the fuzzer; ISG denotes initial seed generation, Inst. denotes instrumentation, SS, ST and SU respectively denote seed selection, seed trim, and seed update (generation or mutation); Feed. denotes feedback mechanism; and BT denotes supported bug types.
Table 1. Characteristics of OS Fuzzers (Sorted by Publication Year)
Trinity(Trinity) Kernel Spec. - - - Gen. Code -
Syzkaller(Syzkaller) Kernel Spec.’ Static Feedback $D$ Mut. Code -
TriforceAFL(TriforceAFL) Kernel Pattern Dynamic Feedback $D$ Mut. Code M.
DIFUZE(2017DIFUZE) Driver Spec.’ Static - - Gen. - M.,L.,O.
VDF(henderson2017vdf) Hypervisor Trace Static - $D$ Mut. Code M.,C.,L.
kAFL(schumilo2017kafl) File System - Static+Dynamic - - Mut. Code M.
SemFuzz(you2017semfuzz) Kernel PoC Static Distance - Mut. Code M.
usb-fuzzer(Syzkaller) Driver Spec.’ Static - - Mut. Code M.
Moonshine(pailoor2018moonshine) Kernel Trace Static Feedback - Mut. Code M.,C.
FUZE(wu2018fuze) Kernel PoC Static - - Mut. Code M.
Schwarz et al.(schwarz2018automated) Kernel Trace - - - Gen. Code C.
Razzer(jeong2019razzer) Kernel Spec.’ Static - - Mut. Code C.
JANUS(JANUS2019fuzzing) File System - Static - - Mut. Code+Cust. M.,L.
SLAKE(chen2019slake) Kernel PoC Static - - Mut. - M.
Shi et al.(shi2019industry) Kernel Spec.’ Static - - Mut. - M.,C.,O.
PeriScope(Song2019PeriScopeAE) Driver Trace Static - - Mut. Code M.
Unicorefuzz(Unicorefuzz2019) Kernel Pattern Static - - Mut. Code M.
KOOBE(chen2020koobe) Kernel PoC Dynamic - - Mut. Code+Cust. M.
Krace(xu2020krace) File System Spec.’ Static Minimum - Mut. Code+Thread C.
Hyper-Cube(schumilo2020hyper) Hypervisor Pattern - - $D$ - - M.,C.,L.,O.
Hydra(Hydra2020finding) File System - Static - - Mut. Code+Cust. M.,L.
USBFuzz(peng2020usbfuzz) Driver Pattern Static - - Mut. Code M.
Agamotto(song2020agamotto) Driver Trace Static - - Mut. Code -
Ex-vivo(Pustogarov2020ExvivoDA) Driver Trace Static - - Mut. Code M.
HFL(kim2020hfl) Kernel Spec.’ Static - - Mut. Code M.
X-AFL(liang2020xafl) Kernel Trace Static - - Mut. Code M.
HEALER(sun2021healer) Kernel Spec.’ Static - - Gen.+Mut. Code M.,C.,L.
Rtkaller(shen2021rtkaller) Kernel Spec.’ Static - - Mut. Code M.,C.
NYX(schumilo2021nyx) Hypervisor Pattern Dynamic - - Mut. Code M.
V-Shuttle(pan2021V-shuttle) Hypervisor Trace Static - - Mut. Code M.,L.
BSOD(maier2021bsod) Driver Pattern Dynamic - - Mut. Code M.
SyzVegas(wang2021syzvegas) Kernel Spec.’ Static Feedback - Mut. Code M.
StateFuzz(zhao2022statefuzz) Driver Spec.’ Static Similarity - Mut. Code+Cust. M.
GREBE(lin2022grebe) Kernel PoC Static Minimum - Mut. Cust. M.,O.
MundoFuzz(myung2022mundofuzz) Hypervisor Trace Static - - Mut. Code M.,L.
Morphuzz(bulekov2022morphuzz) Hypervisor Trace Static - - Mut. Code C.,L.
CONZZER(CONZZER2022context) File System Pattern Static Feedback - Mut. Code+Thread C.
Dr.Fuzz(zhao2022semantic) Driver Spec.’ Static+Dynamic - - Mut. Code+Cust. M.,O.
PrIntFuzz(ma2022printfuzz) Driver Spec.’ Static+Dynamic - - Gen.+Mut. Code M.,C.
DriFuzz(shen2022drifuzz) Driver Trace Static - - Gen.+Mut. Code M.
Hao et al.(Hao2022DemystifyingTD) Kernel - Static - - - - -
KSG(sun2022ksg) Kernel Spec.’ Static - - - - M.,C.,L.
SyzScope(zou2022syzscope) Kernel PoC Static - - Mut. - M.
Tardis(shen2022tardis) Kernel Spec. Static - - Mut. Code M.,L.
SegFuzz(jeong2023segfuzz) Kernel Spec.’ Static - - Mut. Code+Thread M.,C.,O.
IRIS(cesarano2023iris) Hypervisor Trace Static - - Mut. Code M.
FUZZNG(bulekov2023FUZZNG) Kernel Pattern Static - - Mut. Code M.
ACTOR(fleischer2023actor) Kernel Trace Static - - Gen. Code+Cust. M.,L.
SyzDescribe(hao2023syzdescribe) Driver Spec.’ - - - - - -
DEVFUZZ(wu2023devfuzz) Driver Pattern Static+Dynamic - - Mut. Code+Cust. M.,L.,O.
SyzDirect(tan2023syzdirect) Kernel Spec.’ Static Distance - Mut. Code -
ReUSB(Jang2023ReUSB) Driver Trace Dynamic - Protocol Mut. Code M.
DDRace(Yuan2023DDRace) Driver Spec.’ Static Feedback - Mut. Code+Thread M.,C.
VD-Guard(Liu2023VDGuard) Hypervisor Pattern Static - - Mut. Code+Cust. M.,L.
ViDeZZo(Liu2023ViDeZZoDV) Hypervisor Trace Static - - Gen.+Mut. Code M.
Lfuzz(Liu2023LFuzz) File System - Static - - Mut. Code M.,L.
KernelGPT(yang2023kernelgpt) Kernel Spec.’ Static - - Mut. Code M.,O.
BRF (Hung2024BRFFT) Kernel Spec.’ Static - - Gen.+Mut. Code M.,C.
MOCK (Xu2024MOCKOK) Kernel Spec.’ Static - - Gen.+Mut. Code M.,C.
SATURN(Xu2024Saturn) Driver Spec.’ Static Similarity - Gen.+Mut. Code M.,C.,I.
Syzgen++(Chen2024SyzGen++) Driver Spec.’ Static - - - Code M.,C.,I.
VIRTFUZZ(Huster2024ToBoldly) Driver Trace Static - - Mut. Code M.,L.
Monarch(TAO2024MONARCH) File System Spec’ Static - - Mut. Code M.,C.
HYPERPILL(Bulekov2024HYPERPILLFF) Hypervisor Pattern Static - - Mut. Code M.